OpenStudy (anonymous):

WILL FAN AND MEDAL James built a small electric car and recorded the distance it traveled. The table below shows the distance traveled (n) during the first 4 seconds after starting (f). Elapsed Time (seconds) Distance Traveled (feet) 1 6.2 2 12.4 3 18.6 4 24.8

OpenStudy (deving7240):

whats the question?

OpenStudy (anonymous):

Which of the following equations represents the relationship between the distance traveled and the elapsed time?

OpenStudy (anonymous):

Which of the following equations represents the relationship between the distance traveled and the elapsed time? f = 6.2 + n n = 6.2 + f f = 6.2n n = 6.2f

OpenStudy (deving7240):

The fourth one

OpenStudy (anonymous):

thanks